# A Novel Shape-Aware Topological Representation for GPR Data with DNN Integration

### v2

@@ Abstract.
Ground Penetrating Radar (GPR) is a widely used non-destructive testing (NDT) technique for subsurface exploration, particularly in infrastructure inspection and maintenance. However, traditional interpretation methods often suffer from noise sensitivity and limited structural awareness.
This study presents a novel framework that integrates shape-aware topological features—derived from B-scan GPR images using Topological Data Analysis (TDA)—with the object detection capabilities of a YOLOv5-based deep neural network (DNN). The proposed topological representation enhances the geometric salience of buried objects, improving the model’s ability to detect and localize underground utilities, especially pipelines.
To mitigate the scarcity of annotated real-world data, a Sim2Real strategy is employed. Synthetic datasets are generated to simulate diverse subsurface conditions, enabling effective knowledge transfer to real-world scenarios.
Experimental results show consistent improvements in mean Average Precision (mAP), highlighting the robustness and effectiveness of the proposed method. This work demonstrates the potential of TDA-enhanced deep learning for reliable subsurface object detection with broad implications in urban planning, safety inspection, and infrastructure management.
@@
Ground Penetrating Radar (GPR), Topological Data Analysis (TDA), Deep Neural Network (DNN), Object Detection,Sim2Real, Infrastructure Inspection, Pipeline Detection
@@ 1. Introduction
Inspection of subsurface infrastructure–particularly pipelines–is critical for maintaining urban environments and preventing failures [1]. This task necessitates high-accuracy non-destructive testing (NDT) methods [2]. Among these, Ground Penetrating Radar (GPR) has emerged as a widely adopted geophysical technique, facilitating the detection and imaging of subsurface structures across diverse domains such as archaeology, civil engineering, and environmental studies [3, 4]. GPR operates by transmitting electromagnetic waves into the ground and interpreting the resulting reflections to infer subsurface compositions.
Traditionally, the interpretation of GPR signals has relied heavily on manual analysis and physics-based models. These conventional approaches are not only time-consuming and labor-intensive, but also highly susceptible to noise and environmental variability. Such limitations have motivated the exploration of automated solutions. Recent advances in deep learning have driven a paradigm shift in GPR data analysis, enabling significant improvements in automation, detection accuracy, and efficiency. Notably, object detection models–including YOLO (You Only Look Once) [5], Faster R-CNN [6], and U-Net [7]–have demonstrated considerable promise. YOLO excels in rapid classification of subsurface anomalies (e.g., pipelines, voids), requiring minimal post-processing. In contrast, Faster R-CNN often achieves higher accuracy, especially for small or low-contrast objects, but at the cost of increased computational demands, though it typically incurs greater computational cost [8]. U-Net is particularly effective for pixel-level segmentation, aiding in the delineation of complex subsurface structures, though it is prone to overfitting when annotated data is scarce.
@@
Despite these advances, a key limitation persists: the performance of deep learning models is heavily reliant on the availability and diversity of high-quality annotated data. Collecting labeled GPR datasets is expensive and logistically challenging, prompting increased interest in simulation-based training approaches.
Sim-to-Real (Sim2Real) transfer has emerged as a viable solution by enabling models trained on simulated GPR data to generalize to real-world environments [14, 15]. Simulated environments allow for the generation of diverse, labeled, and noise-controlled data representing various subsurface conditions. These approaches help bridge the domain gap between synthetic and field data, improving model generalization and robustness.
@@
In this paper, we present a deep learning framework that incorporates both Sim2Real training and TDA-based feature augmentation within a deep learning YOLOv5 pipeline. Our method demonstrates improved detection performance, as measured by mean Average Precision(mAP), and offers a robust, scalable solution for subsurface object detection.
The remainder of this paper is organized as follows. Section 2 introduces the fundamental concepts of persistent homology that underpin our approach. Section 3 details the proposed framework, including methodology and implementation. Section 4 presents the experimental setup and reports the results on selected datasets.
2. Persistent Homology
@@
3. Methodology
3.1. Step 1: Shape-Aware Topological Representation
@@
We treat each B-scan image as a grayscale 2D image and apply the persistent homology framework described in Section 2. We focus exclusively on $\beta_{1}$ topological features (loops), as our primary interest lies in the crescent-shaped closed patterns representing subsurface pipelines.
The key idea of our approach lies in transforming abstract topological information into a format that CNNs can effectively utilize. Rather than relying on conventional persistence diagrams or barcodes, which provide compact summaries but cannot be directly processed by CNNs, we construct a spatial representation that preserves both topological significance and spatial locality. We create a new image with identical dimensions to the original input, visualizing the shapes of degree-one homology generators while modulating their intensity according to their lifetime. Generators with longer lifetimes are rendered with higher intensity, while shorter-lived features appear with lighter shading. This spatial encoding enables CNNs to simultaneously access local pixel information and global topological structure within their standard convolution operations. To construct the final input for the object detection model, we stack the grayscale GPR B-scan image and the lifetime-weighted topological feature map along the channel dimension, forming a multi-channel representation. This approach allows the CNN-based detector (YOLOv5) to learn from both raw spatial information and topological cues, effectively augmenting its shape perception capabilities. The CNN can thus leverage its existing architectural strengths for local feature detection while gaining access to global topological information that would otherwise be inaccessible through standard convolution operations.
@@
This represents a novel contribution that enables CNNs to overcome their inherent shape-blindness by providing explicit topological guidance encoded in a spatially coherent manner. To the best of our knowledge, this approach of encoding topological lifetime information as spatial intensity maps for CNN consumption has not been addressed in prior research. An example of such a visualization, constructed from real GPR B-scan data using the proposed method, is shown in Figure 2.
@@ 3.2. Step 2: Sim2Real Training
To overcome this data scarcity challenge, we adopt a Sim2Real training philosophy that leverages synthetic data generation followed by real-world fine-tuning. Our approach begins with pre-training the model on a comprehensive dataset of synthetic GPR B-scan images generated using gprMax, with the data generation procedure detailed in Section IV. 4.1. These simulated images undergo processing through our TDA pipeline (Section 3.1), where $\beta_{1}$ topological features corresponding to hyperbolic reflections are extracted and encoded into lifetime-weighted shape-aware representations.
@@
This two-stage approach maximizes the utility of limited real-world data while leveraging the scalability of synthetic data generation, ultimately achieving robust performance in practical deployment scenarios. The complete training schema is illustrated in Figure 3.
@@ 4.1. Numerical Data Generation
The simulation setup included essential components such as pipe placement, soil layering, and grid resolution to reflect realistic underground environments. A Ricker wavelet–a zero-phase wavelet commonly used in GPR applications due to its compact time-domain representation–with a 350MHz center frequency was employed as the excitation signal. The radar system was modelled with a co-located transmitter-receiver pair moved in 0.024m increments along the surface, resulting in a total of 456 A-scans (single-point radar measurements) per simulation to construct each B-scan cross-sectional image.
@@
B-scan images were generated through 2D electromagnetic simulations under both homogeneous and heterogeneous conditions, with scenarios with single and multiple pipes. As shown in Figure 4, wave reflections vary depending on the complexity of subsurface configuration. To improve visibility of deep-layer structures, additional signal post-processing techniques were applied. To further improve the quality of the simulated GPR signals, we applied several signal processing techniques, including background removal (to eliminate static noise), band-pass filtering (to isolate the relevant frequency range), and Automatic Gain Control (AGC) for contrast normalization. The band-pass filter was applied in the frequency range of $100~{}\mathrm{MHz}$ to $1900~{}\mathrm{MHz}$ to retain informative signal components while suppressing irrelevant noise. For each B-scan image, five AGC variants were generated to increase the model’s robustness against intensity variations. In total, 1,200 simulated B-scan images were produced. An example output is shown in Figure 4(a).
@@ 4.2. Field Data Collection
Figure 4. Examples of B-scan images used in training. (a) Simulated sample data generated from a synthetic underground model. (b) Field data collected from a controlled testbed environment.
Post-processing was performed using Radan 7 software. Standard GPR signal processing steps were applied, including background subtraction, band-pass filtering centered around the antenna frequency, and time-varying gain adjustments. These procedures were consistent with those used for simulated data, ensuring a uniform preprocessing pipeline across domains.
@@
Table 2. Descriptive statistics of B-scan datasets
@@
Simulated 1080×1080 300/300 280$\pm$33 × 167$\pm$43
Site A 680×680 86/94 390$\pm$87 × 107$\pm$23
Site B 680×680 70/70 234$\pm$53 × 79$\pm$19
Site C 680×680 11/13 323$\pm$134 × 129$\pm$38
4.3. Model Configurations
(2) $F_{\mathrm{val}}=F^{(*)},\quad F_{\mathrm{tr}}=\bigcup_{j\neq*}F^{(j)},$
[…]
Table 3. Model configurations and dataset composition
Model Training Data Validation Data Test Data

### v3

# Shape-Aware Topological Representation for Pipeline Hyperbola Detection in GPR Data

@@ Abstract.
Ground Penetrating Radar (GPR) is a widely used non-destructive testing (NDT) technique for subsurface exploration, particularly in infrastructure inspection. However, traditional interpretation methods often struggle with noise sensitivity and limited structural awareness. We propose a novel framework that integrates shape-aware topological features, derived from B-scan GPR images using Topological Data Analysis (TDA), with the object detection capabilities of a YOLOv5-based deep neural network (DNN). This topological representation improves geometric salience, enhancing detection and localization of underground utilities, especially pipelines. To mitigate the scarcity of annotated real-world data, a Sim2Real strategy is employed. Synthetic datasets are generated to capture both diverse subsurface conditions and the essential hyperbolic reflection patterns of pipelines, enabling more effective knowledge transfer to real-world scenarios. Experimental results show consistent improvements in mean Average Precision (mAP), highlighting the robustness and effectiveness of the proposed method. This work demonstrates the potential of TDA-enhanced deep learning for reliable subsurface object detection with broad implications in urban planning, safety inspection, and infrastructure management.
@@
Ground Penetrating Radar (GPR), Topological Data Analysis (TDA), Shape-aware representation, Sim2Real transfer, Subsurface object detection
@@ 1. Introduction
Inspection of subsurface infrastructure–particularly pipelines–is critical for maintaining urban environments and preventing failures [1]. This task necessitates high-accuracy non-destructive testing (NDT) methods [2]. Among these, Ground Penetrating Radar (GPR) has emerged as a widely adopted geophysical technique, facilitating the detection and imaging of subsurface structures across diverse domains such as archaeology, civil engineering, and environmental studies [3, 4]. GPR operates by transmitting electromagnetic waves into the ground and interpreting the resulting reflections to infer subsurface compositions. Traditionally, the interpretation of GPR signals has relied heavily on manual analysis and physics-based models. These conventional approaches are not only time-consuming and labor-intensive, but also highly susceptible to noise and environmental variability. Such limitations have motivated the exploration of automated solutions.
Recent advances in deep learning have driven a paradigm shift in subsurface sensing tasks [5, 6, 7, 8, 9]. Similarly, in Ground Penetrating Radar (GPR), deep learning techniques have been applied to enable significant improvements in automation, detection accuracy, and computational efficiency. Notably, object detection models–including YOLO (You Only Look Once) [10], Faster R-CNN [11], and U-Net [12]–have demonstrated considerable promise. YOLO excels in rapid classification of subsurface anomalies (e.g., pipelines, voids), requiring minimal post-processing. In contrast, Faster R-CNN often achieves higher accuracy, especially for small or low-contrast objects, but at the cost of increased computational demands, though it typically incurs greater computational cost [13]. U-Net is particularly effective for pixel-level segmentation, aiding in the delineation of complex subsurface structures, though it is prone to overfitting when annotated data is scarce.
@@
Despite these advances, a key limitation persists: the performance of deep learning models is heavily reliant on the availability and diversity of high-quality annotated data. Collecting labeled GPR datasets is expensive and logistically challenging, as field measurements are often limited by access, cost, and environmental constraints. Recent studies have explicitly highlighted the scarcity of publicly available GPR datasets for underground utility detection, emphasizing that the lack of large-scale, standardized field data remains a critical bottleneck for deep-learning-based interpretation [19, 20]. This motivates the increased adoption of simulation-based training and domain adaptation strategies.
Sim-to-Real (Sim2Real) transfer has emerged as a viable solution by enabling models trained on simulated GPR data to generalize to real-world environments [21]. Simulated environments allow for the generation of diverse, labeled, and noise-controlled data representing various subsurface conditions. These approaches help bridge the domain gap between synthetic and field data, improving model generalization and robustness.
@@
Unlike conventional TDA–CNN approaches that employ statistical embeddings such as persistence images (PI) [27], persistence entropy (PE) [28], or persistence landscapes (PL) [29], which summarize persistence diagrams into abstract feature vectors, our method preserves spatial coherence through a shape-aware topological representation. This representation directly encodes the geometric continuity and connectivity of reflection hyperbolas within the image domain, allowing topological information to be physically aligned with pixel-level GPR reflections. Consequently, it bridges the gap between topological abstraction and geometric interpretability, providing a more intuitive and robust integration of TDA into deep-learning-based GPR analysis.
In this paper, we propose a novel framework, TE-S2R (TDA-Enhanced Sim2Real), for robust subsurface object detection. The core idea is to augment conventional deep learning pipelines with shape-aware topological features extracted via TDA, which effectively capture structural information while remaining resilient to noise. Unlike prior work, our approach directly incorporates topological features as enhanced image representations, enabling the model to learn both geometric and spatial characteristics more effectively.
A key challenge in Sim2Real approaches lies not in reducing data collection costs, but in bridging the distributional gap between synthetic and real-world data, particularly with respect to structural inconsistencies and noise. TE-S2R addresses this issue by integrating TDA-based features that emphasize shape and connectivity over raw pixel values. This reduces sensitivity to domain-specific noise and enhances topological alignment between synthetic and real GPR images. As a result, our method facilitates more stable and reliable knowledge transfer across domains. The synergy between TDA and Sim2Real yields a significant performance boost under real-world conditions, where noise and variability are unavoidable.
The resulting TE-S2R framework demonstrates consistent improvements in mean Average Precision (mAP), offering a scalable and noise-resistant solution that is effective not only for GPR data, but also extensible to other types of grid-based imaging modalities.
2. Methodology
2.1. Step 1: Shape-Aware Topological Representation
@@
We treat each B-scan image as a grayscale 2D image and apply the persistent homology framework. We focus exclusively on $\beta_{1}$ topological features (loops), as our primary interest lies in the crescent-shaped closed patterns representing subsurface pipelines.
The key idea of our approach lies in transforming abstract topological information into a format that CNNs can effectively utilize. Rather than relying on conventional persistence diagrams or barcodes, which provide compact summaries but cannot be directly processed by CNNs, we construct a spatial representation that preserves both topological significance and spatial locality. We create a new image with identical dimensions to the original input, visualizing the shapes of degree-one homology generators while modulating their intensity according to their lifetime. Generators with longer lifetimes are rendered with higher intensity, while shorter-lived features appear with lighter shading.
The detailed procedure for constructing the lifetime-weighted topological feature map is summarized in Algorithm 1. Each GPR B-scan is first converted into a grayscale matrix and processed through a cubical filtration to compute persistent homology using the Ripserer.jl package implemented in the Julia environment. The resulting $H_{1}$ generators representing loop-like topological structures are extracted along with their birth and death times, and their normalized lifetimes $(d_{i}-b_{i})$ are used as transparency weights $\alpha_{i}$ to indicate topological significance. These generators are then visualized on an image grid and their weighted intensities are accumulated onto the original B-scan domain, producing a spatially aligned and topologically enhanced representation $T$ that preserves both amplitude and geometric stability. This fused map serves as the final input to the CNN backbone without requiring any architectural modification.
Algorithm 1 Generation of Lifetime-Weighted Topological Feature Map
1:Input: B-scan image II
2:Output: Lifetime-weighted feature map $T$
3: Convert II to grayscale matrix $G\in[0,1]$
4: Reverse $G$ vertically to align coordinate orientation
5: Compute persistent homology using cubical filtration: $PH\leftarrow\texttt{ripserer(Cubical($G$); cutoff=0.2)}$
6: Extract $H_{1}$ generators with representative cycles from PHPH
7:if $H_{1}\neq\emptyset$ then
8:  for each generator $g_{i}\in H_{1}$ do
9:   Compute lifetime $l_{i}=d_{i}-b_{i}$
10:  end for
11:  Normalize $l_{i}$ to $[0.3,1.0]$ as transparency weights $\alpha_{i}$
12:  Initialize $T\leftarrow\mathbf{0}$ with the same size as $G$
13:  for each generator $g_{i}\in H_{1}$ do
[…]
15:    Draw a line between $(v_{1},v_{2})$ on $H$ with color = green, opacity = $\alpha_{i}$
16:    Accumulate $\alpha_{i}$ into $T$ at corresponding coordinates
17:   end for
18:  end for
19:else
20:  Skip visualization (no $H_{1}$ features detected)
21:end if
22: Save the overlay visualization $H$ and feature map $T$
This spatial encoding enables CNNs to simultaneously access local pixel information and global topological structure within their standard convolution operations. To construct the final input for the object detection model, we stack the grayscale GPR B-scan image and the lifetime-weighted topological feature map along the channel dimension, forming a multi-channel representation. This approach allows the CNN-based detector (YOLOv5) to learn from both raw spatial information and topological cues, effectively augmenting its shape perception capabilities. The CNN can thus leverage its existing architectural strengths for local feature detection while gaining access to global topological information that would otherwise be inaccessible through standard convolution operations.
@@
This represents a novel contribution that enables CNNs to overcome their inherent shape-blindness by providing explicit topological guidance encoded in a spatially coherent manner. To the best of our knowledge, this approach of encoding topological lifetime information as spatial intensity maps for CNN consumption has not been addressed in prior research. An example of such a visualization, constructed from real GPR B-scan data using the proposed method, is shown in Fig. 2.
@@ 2.2. Step 2: Sim2Real Training
To overcome this data scarcity challenge, we adopt a Sim2Real training philosophy that leverages synthetic data generation followed by real-world fine-tuning. Our approach begins with pre-training the model on a comprehensive dataset of synthetic GPR B-scan images generated using gprMax, with the data generation procedure detailed in Section IV. 3.1. These simulated images undergo processing through our TDA pipeline (Section 2.1), where $\beta_{1}$ topological features corresponding to hyperbolic reflections are extracted and encoded into lifetime-weighted shape-aware representations.
@@
This two-stage approach maximizes the utility of limited real-world data while leveraging the scalability of synthetic data generation, ultimately achieving robust performance in practical deployment scenarios. The complete training schema is illustrated in Fig.  3.
@@ 3.1. Numerical Data Generation
The simulation setup included essential components such as pipe placement, soil layering, and grid resolution to reflect realistic underground environments. A Ricker wavelet–a zero-phase wavelet commonly used in GPR applications due to its compact time-domain representation–with a 350MHz center frequency was employed as the excitation signal. The radar system was modelled with a co-located transmitter–receiver pair that moved along the surface to acquire successive A-scans forming each B-scan cross-sectional image.
@@
The specific input parameters used for the Peplinski and fractal models are summarized in Table 1. These include the sand and clay content, bulk and particle density, and the range of volumetric water content, as well as the fractal dimension ($D_{\text{frac}}$) and directional weights ($W_{\text{frac}}$). These parameters were selected within realistic ranges to generate diverse dielectric and conductive profiles for pretraining the Sim2Real transfer-learning model, rather than to reproduce site-specific soil conditions. This approach follows the common practice in GPR simulation studies, where synthetic data are primarily used to initialize deep-learning models before real-data adaptation [36].
The 2D simulation setup is illustrated in Fig.  4(a). The transmitter and receiver antennas were positioned 12 cm apart and moved together along the survey direction in 0.024 m increments per A-scan, resulting in 456 traces per B-scan. Both antennas were placed directly on the ground surface without a separation gap, while the air layer above the surface was modeled with standard air properties.
Table 1. Peplinski and Fractal Input Parameters Used in the FDTD Simulation
Model $S$ (%) $C$ (%) $\rho_{b}$ (g/cm3) $\rho_{s}$ (g/cm3) $w_{\text{range}}$
Peplinski 80 20 2.0 2.66 [0.01–0.10]
Fractal $N_{\text{mat}}$ $D_{\text{frac}}$ $W_{\text{frac}}$ [x, y, z]
10 1.0 [2.0, 0.1, 0.1]
@@
B-scan images were generated through 2D electromagnetic simulations under both homogeneous and heterogeneous conditions, with scenarios with single and multiple pipes.To improve the quality and interpretability of the synthetic B-scans, standard GPR preprocessing operations (background removal, band-pass filtering, and automatic gain control) were applied, consistent with the procedures used for real data (see Section 3.2).
For each B-scan image, five AGC variants were generated to increase the model’s robustness against intensity variations. In total, 300 simulated B-scan images were produced. An example output is shown in Fig.  4(c).
It is important to note that the simulated data were not generated to replicate any specific field site. Instead, the simulation stage was designed to facilitate the learning of fundamental GPR reflection patterns (e.g., hyperbolic responses) under controlled conditions. The real-world datasets (Sites A, B, and C) were independently collected from distinct geographical areas with different soil types and noise characteristics. Hence, the proposed Sim2Real framework focuses on transferring generalizable feature representations from synthetic to heterogeneous real-world environments, rather than modeling a single site-specific scenario.
@@ 3.2. Field Data Collection
According to the manufacturer’s documentation [37, 39, 38], the 350 MHz 350HS antenna provides a typical detection depth of approximately 6 m and a maximum penetration of up to 12 m, depending on soil moisture and dielectric properties. The effective interpretation depth was empirically determined from the signal-to-noise ratio and the visibility of reflection hyperbolas in the processed B-scan images. In the collected field data, clear hyperbolic reflections were consistently observed up to about 3 m below the surface, beyond which the signals became indistinct due to increased attenuation in moist and conductive soils, indicating that the radar performance was suitable for the experimental conditions of this study.
Figure 4. Overview of the experimental setup and representative data samples. (a) Schematic diagram of the simulation model showing the pipe geometry, grid configuration, and boundary conditions, with relevant dimensions labeled (pipe diameter = 0.5 m, burial depths = 0.3–0.4 m, grid resolution = 6 mm). (b) On-site photograph of the field data collection setup showing the GPR scanning apparatus and buried pipe layout. (c) Example of a simulated B-scan image generated from the synthetic model corresponding to the setup in (a). (d) Example of a real B-scan image collected from Site A, corresponding to a 2 m survey line and a buried pipe at a depth of approximately 0.5 m.
Both simulated and field B-scan datasets were processed using the same preprocessing pipeline to ensure consistent data quality. The procedures included background removal, automatic gain control (AGC), and band-pass filtering. Background removal was implemented by subtracting the median of 11 consecutive rows to eliminate static noise components. AGC was applied with a moving window of 151 rows and gain factors $G_{\mathrm{max}}=[1.0,10,10^{2},10^{3},10^{4},10^{5}]$ to enhance deeper reflections. A second-order band-pass filter (100–1500 MHz) was used to suppress low- and high-frequency noise. Table 3 summarizes the input parameters used in the preprocessing steps. For simulated data, 1024 uniformly spaced time samples were selected to match the resolution of the field data.
Table 3. Input parameters applied in signal processing.
Process Parameter Symbol Value
Background removal Window size (row) $w_{\mathrm{bg}}$ 11
Auto Gain Control Gain factor $g_{\max}$ [1.0, 10, 102,
103, 104, 105]
Window size (row) $w_{\mathrm{AGC}}$ 151
Band-pass filter High/Low cutoff (MHz) $(f_{H},f_{L})$ (1500, 100)
Filter order 2
@@
Although the number of real-world samples is limited (167 images in total, with 11 reserved for independent testing), statistical reliability was ensured using a five-fold cross-validation scheme across Sites A and B. In addition, multiple data augmentation techniques were applied in the YOLOv5 model to enhance sample diversity and prevent overfitting.
Although the simulated and real datasets have different original resolutions (1080×1080 and 680×680, respectively) and sample sizes (300 synthetic vs. 167 field images), all B-scan images were resized to 416×416 and underwent identical preprocessing and augmentation steps before training. This ensures consistent input dimensions and comparable learning conditions across all configurations. Moreover, the simulated data were used solely for pretraining to learn generalizable reflection patterns, while the real datasets were used for fine-tuning and evaluation to assess domain transferability.
It should be noted that the simulated data were not generated to replicate any specific field site. Instead, the simulation stage was designed to provide a controlled environment for learning fundamental GPR signal features—such as hyperbolic responses and subsurface reflections—under idealized conditions. The real-world datasets used for training and testing were independently collected from different geographical locations (Sites A, B, and C), each exhibiting distinct soil compositions and noise characteristics. Therefore, the proposed Sim2Real framework aims to transfer generalizable feature representations learned from simulation to heterogeneous real-world domains, rather than modeling a single site-specific scenario.
Table 4. Descriptive statistics of B-scan datasets.
@@
Simulated $1080\times 1080$ 300/300 $280\!\pm\!33~\times~167\!\pm\!43$
Site A $680\times 680$ 86/94 $390\!\pm\!87~\times~107\!\pm\!23$
Site B $680\times 680$ 70/70 $234\!\pm\!53~\times~79\!\pm\!19$
Site C $680\times 680$ 11/13 $323\!\pm\!134~\times~129\!\pm\!38$
Table 5. Model configurations and dataset composition. Training and validation subsets are randomly divided from the combined data of Sites A and B, while testing is performed exclusively on independent Site C.
Model Train (Sim, A+B) Validation (A+B) Test (Site C)
@@
Note: The simulated dataset contains 300 samples. The field dataset includes 156 samples (125 for training, 31 for validation per fold), and the test set consists of 11 independent samples from Site C.
3.3. Model Configurations
(2) $F_{\mathrm{val}}=F^{(*)},\quad F_{\mathrm{tr}}=\bigcup_{j\neq*}F^{(j)},$
where $\in\{1,2,3,4,5\}$ denotes the fold index for the current validation set. These notations and data splits are consistently applied across all configurations to ensure fair comparison. In addition, 11 real-world samples from Site C are employed as test dataset.
For all model configurations, the dataset is divided into three disjoint subsets—training, validation, and testing.
  • Model Base and Model TE-Base, the training and validation subsets are drawn from the combined field data of Sites A and B, comprising 125 samples for training and 31 samples for validation in each fold, following a five-fold cross-validation scheme to ensure statistical robustness. The folds were stratified such that samples from both sites were evenly distributed across all folds, thereby avoiding any site-specific bias in training or validation.
  • Model S2R and Model TE-S2R are first pre-trained using 300 simulated B-scan images and subsequently fine-tuned with the same training and validation protocol as Model Base and Model TE-Base.
  • Model Sim and Model TE-Sim are trained solely on the 300 simulated B-scans but evaluated using the same five-fold validation process for consistency.
All configurations are finally tested on the independent Site C dataset (11 samples), which is completely excluded from training and fine-tuning, providing a strict evaluation of generalization performance under domain-shift conditions.